# Mixture of Group Experts for Learning Invariant Representations

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06submissionid: 4041
@@ 1. Introduction
Despite their advantages, MoE models face a critical challenge: the lack of diversity and specialization among experts (Zhou et al., 2022; Sun et al., 2024; Wang et al., 2024; Feng et al., 2025; Nakamura et al., 2025). In an MoE layer, a gating network determines token-to-expert mappings, ideally routing similar or relevant tokens to the same group of experts. However, the vanilla MoE layer often produces redundant experts that fail to achieve sufficient diversity and specialization, thereby undermining the performance of MoE models. Promoting diversity and specialization among experts remains a significant challenge for MoE models, especially as the number of experts scales.
@@
2. Related Work
2.1. Mixture-of-Experts
Mixture-of-Experts (MoE) was first introduced in the early 1990s (Jacobs et al., 1991; Jordan and Jacobs, 1994). Later, Shazeer et al. (Shazeer et al., 2017) proposed integrating an MoE layer between LSTM layers, leading to remarkable advancements in machine translation. This idea was further extended by incorporating the MoE layer into the Transformer architecture, replacing traditional feedforward layers. For example, GShard (Lepikhin et al., 2021) refines the top-2 routing strategy, significantly boosting machine translation performance. Switch Transformer (Fedus et al., 2022) simplifies the gating mechanism by selecting only the top-1 expert per token, achieving superior scalability. More recently, PEER (He, 2024) introduces an expansion of the expert count in MoE to one million, showcasing its potential in pushing the boundaries of MoE models. Commercial-scale language models have also adopted MoE architectures, with notable examples including Mixtral-8x7B (Jiang et al., 2024), Qwen2.5 (Yang et al., 2024b), and DeepSeek-V3 (Liu et al., 2024).
2.2. Promoting Expert Diversity
Researchers have developed a variety of strategies to construct diverse experts, drawing inspiration from ensemble learning (Wood et al., 2023). Some methods focus on leveraging different initialization techniques. For instance, MMoE (Ma et al., 2018) achieves expert diversity by introducing varying degrees of randomness into both the training data and model initialization processes. Similarly, MoE Jetpack (Zhu et al., 2024) employs diverse dense checkpoints to initialize different experts, while Drop-Upcycling (Nakamura et al., 2025) takes an alternative approach by using pre-trained dense models for initialization, followed by a strategic re-initialization of specific weight components to enhance diversity. Other approaches aim to diversify experts by training them on distinct datasets. For example, MoCLE (Gou et al., 2023) clusters the training data into multiple groups using a pre-trained clustering model and learns to activate specific LoRA experts at each layer based on the data cluster. Structural variations have also proven effective for promoting expert diversity. MLoRE (Yang et al., 2024a) utilizes low-rank experts with varying ranks, while HMoE (Wang et al., 2024) and MoDSE (Sun et al., 2024) achieve diversity by incorporating experts of different sizes. Advances in token-to-expert mapping represent another avenue for innovation. MMoEEx (Aoki et al., 2022), for instance, employs exclusion and exclusivity mechanisms to induce diversity among experts. Meanwhile, Zhou et al. (Zhou et al., 2022) propose an alternative approach through expert-choice routing, wherein tokens may be processed by a variable number of experts rather than adhering to conventional token-choice routing mechanisms. Additionally, shared experts have garnered attention in recent works (Han et al., 2024; Liu et al., 2024), demonstrating significant adaptability and efficiency by remaining universally activated. Regularization techniques further enhance expert diversity. For example, MOORE (Hendawy et al., 2024) combines the Gram-Schmidt process with reinforcement learning to orthogonalize expert-generated representations, promoting decorrelation among experts. Similarly, OMoE (Feng et al., 2025) applies an orthogonalization technique during fine-tuning. Meanwhile, MoELoRA (Luo et al., 2024) employs a contrastive learning framework, treating outputs from the same expert as positive samples and those from different experts as negative ones.
@@ 3. Methodology
In this section, we begin by introducing the vanilla MoE, followed by an analysis of MoE from the perspective of sparse representation. We then provide a comprehensive explanation of the implementation process and underlying mechanism of our proposed MoGE. Finally, we investigate how MoGE effectively learns invariant representations.
@@ 3.1. Vanilla MoE
For simplicity, we omit considerations of noise, load balancing loss, and any sequence dependencies between the softmax and the top-$k$ functions. In our experiments, we follow the default settings for each MoE model outlined in its original paper.
@@ 3.2. Sparse Representation for MoE
3.3. Mixture of Group Experts
Figure 2. Illustrating the computation process of our group sparse regularization.
@@ 3.3. Mixture of Group Experts
We first introduce the computation process of our group sparse regularization for $z$ and then provide an analysis of the underlying reasoning. The computation process is illustrated in Fig. 2. Given a vector $z\in\mathbb{R}^{n}$, we begin by reshaping it into an approximately square matrix $Z\in\mathbb{R}^{r\times c}$, where $n=r\times c$ and $r$ is chosen to be as close to $c$ as possible. Next, we compute the elementwise square of $Z$ to obtain $\tilde{Z}$, i.e., $\tilde{Z}=Z\odot Z$, where $\odot$ denotes the elementwise product. Subsequently, we generate a rotationally symmetric Gaussian lowpass filter $F_{\sigma}$ of size $h\times h$ with a standard deviation $\sigma$. We then convolve $\tilde{Z}$ with the Gaussian filter $F_{\sigma}$ to produce $\tilde{P}$, i.e., $\tilde{P}=F_{\sigma}*\tilde{Z}$, where * denotes the 2D convolution. Finally, we compute the square root of each element in $\tilde{P}$ to obtain $P$. The group sparse regularization $R(z)$ is defined as the sum of all elements in $P$. We summarize the entire computation process of our group sparse regularization in Algorithm 1.
Table 1. Comparison of the average image Euclidean distances in $Z$ between the original images from the Fashion-MNIST test set and their corresponding variations under different transformations. A smaller value indicates a higher degree of invariance in $Z$ to the respective transformation.
Type Rotation Scaling Translation Shear
@@
where $\sigma_{0}>0$ and $\sigma_{\min}>0$ are constants satisfying $\sigma_{0}\geq\sigma_{\min}$, $\gamma\geq 0$, $t$ represents the current iteration number, and $T$ is the total number of iterations. The curves of $\sigma$ with varying $\gamma$ over iterations are illustrated in Appendix A. Finding the optimal scheduler for $\sigma$ is a challenging task. This complexity arises not only from the intricate interplay between learning rate, batch size, number of experts $n$, the $k$ value, and other hyperparameters but also from the prohibitively high cost of conducting a hyperparameter search for MoE models with massive parameters (Bergsma et al., 2025). We will explore optimal scheduler for $\sigma$ in the future.
@@
where $\lambda>0$ balances the original loss function and the regularization. Our group sparse regularization requires computing a 2D convolution with a Gaussian lowpass filter (see Algorithm 1). In practice, we empirically use valid convolution to avoid padding operations on $\tilde{Z}$ (see Fig. 2). Additionally, we need to predefine the filter size $h$ and the standard deviation $\sigma$. In subection 4.4, we will conduct ablation experiments to analyze these aspects.
@@ 3.4. Investigating Invariant Representations
We explore how MoGE learns invariant representations by employing a simplified model with a single MoE layer for image classification on the Fashion-MNIST dataset (Xiao et al., 2017). The model consists of 400 homogeneous experts with top-1 routing, where each image is treated as a token. Each expert is a two-layer feedforward network, featuring 784 neurons in both the input and output layers, and 64 neurons in the hidden layer. The output of the MoE layer is fed directly into a single-layer feedforward classifier for categorization into 10 classes.
We train MoE and MoGE models from scratch for 150 epochs, setting $h=3$, $\sigma=2$, and $\lambda=$ 4e-3 for MoGE. The top-1 test accuracies achieved by MoE and MoGE are $41.70\%$ and $44.74\%$, respectively. Due to the large number of experts and the relatively small size of the training set, these test accuracies are not particularly high. However, our primary focus here is on the relative performance of the two models. Fig. 3 illustrates the average $Z$ values in Algorithm 1 for both MoE and MoGE across each category in the Fashion-MNIST test set. Notably, the average elements of $Z$ in MoGE exhibit a clear aggregation effect throughout the training process. In contrast, this aggregation effect is not observed in MoE. This indicates that the experts activated for each category in MoGE exhibit spatial similarity, which helps enhance the diversity and specialization of the experts.
To validate that MoGE effectively learns invariant representations, we apply various transformations to the images in the Fashion-MNIST test set and evaluate the image Euclidean distances (IMED) (Wang et al., 2005) in $Z$ between the original and the transformed images. Unlike the standard Euclidean distance, which assumes spatial independence of pixels and may produce counterintuitive results (e.g., assigning smaller distances to perceptually significant distortions), IMED takes into account the spatial correlations between pixels. This characteristic makes IMED a widely adopted metric for comparing 2D images. As shown in Table 1, we use the default function RandomAffine in PyTorch to implement four transformations: rotation, scaling, translation, and shear. The rotation transformation rotates input images by angles of 5, 10, and 15 degrees. The scaling transformation applies scale factors of 0.5, 0.8, and 1.1. The translation transformation shifts images horizontally and vertically by (0, 0.1), (0.1, 0), and (0.1, 0.1), relative to the image size. The shear transformation utilizes shear angles of 5, 10, and 15 degrees. The results reported in Table 1 demonstrate that following these transformations, the average IMED values in $Z$ for MoGE are consistently smaller than those for MoE. This highlights the significant advantage of MoGE in learning more invariant representations.
@@ 4. Experiments
In this section, we first evaluate the superiority of MoGE on image classification and language modeling tasks. Subsequently, we test the scalability and efficiency of MoGE, followed by ablation studies on the Gaussian lowpass filter.
Table 2. Comparison of the dense, MoE, and our MoGE models for training from scratch and fine-tuning in image classification.
Dataset Model Resolution Acc@1 Acc@5 $n$ $k$Params #Paramsact
CIFAR-100 ViT-S/4 (Dosovitskiy et al., 2020) 74.31 91.34 - - 25.3M 25.3M
V-MoE-S/4 (Riquelme et al., 2021) $32\times 32$ 74.03 91.49 32 1 155.5M 25.4M
V-MoGE-S/4 74.67 91.55 32 1 155.5M 25.4M
SwinV2-T/4 (Liu et al., 2022a) 76.62 92.09 - - 27.7M 27.7M
SwinV2-MoE-T/4 (Hwang et al., 2023) $32\times 32$ 77.35 92.41 32 1 284.0M 27.7M
SwinV2-MoGE-T/4 77.75 92.61 32 1 284.0M 27.7M
Tiny-ImageNet ViT-S/8 (Dosovitskiy et al., 2020) 55.11 76.25 - - 25.5M 25.5M
V-MoE-S/8 (Riquelme et al., 2021) $64\times 64$ 55.38 76.66 32 1 155.7M 25.5M
V-MoGE-S/8 56.15 77.24 32 1 155.7M 25.5M
SwinV2-T/7 (Liu et al., 2022a) 58.42 77.72 - - 27.7M 27.7M
SwinV2-MoE-T/7 (Hwang et al., 2023) $64\times 64$ 59.01 78.97 32 1 284.1M 27.8M
SwinV2-MoGE-T/7 60.19 79.50 32 1 284.1M 27.8M
ImageNet-1K ViT-S/16 (Dosovitskiy et al., 2020) 77.81 93.78 - - 26.2M 26.2M
V-MoE-S/16 (Riquelme et al., 2021) $224\times 224$ 76.95 92.95 32 1 156.4M 26.3M
V-MoGE-S/16 77.24 93.18 32 1 156.4M 26.3M
SwinV2-T/12 (Liu et al., 2022a) 80.43 95.30 - - 28.3M 28.3M
SwinV2-MoE-T/12 (Hwang et al., 2023) $192\times 192$ 79.14 94.00 32 1 211.4M 28.4M
SwinV2-MoGE-T/12 79.97 94.94 32 1 211.4M 28.4M
ViT-S/16 (Dosovitskiy et al., 2020) 81.47 96.20 - - 22.1M 22.1M
V-MoE-S/16 (Riquelme et al., 2021) $224\times 224$ 81.19 96.07 32 1 95.3M 22.1M
ImageNet-1K V-MoGE-S/16 81.50 96.19 32 1 95.3M 22.1M
(fine-tuning) SwinV1-T/7 (Liu et al., 2021) 80.80 95.94 - - 28.3M 28.3M
SwinV1-MoE-T/7 (Hwang et al., 2023) $224\times 224$ 80.81 95.86 32 1 211.3M 28.3M
SwinV1-MoGE-T/7 81.06 95.92 32 1 211.3M 28.3M
@@ 4.1. Image Classification
Setup. We leverage three commonly used datasets: CIFAR-100 (Krizhevsky and Hinton, 2009), Tiny-ImageNet (Le and Yang, 2015), and ImageNet-1K (Russakovsky et al., 2015). For CIFAR-100 and Tiny-ImageNet, we apply data augmentation techniques as described in (Lee et al., 2021), while adhering to standard practices (Liu et al., 2022a, 2021; Wang et al., 2021; Liu et al., 2022b) for ImageNet-1K. We include three widely recognized Transformer models for comparison: Vision Transformer (ViT) (Dosovitskiy et al., 2020), Swin Transformer V1 (SwinV1) (Liu et al., 2021), and Swin Transformer V2 (SwinV2) (Liu et al., 2022a).
@@
Consistent with prior studies, the architecture of each expert within the MoE layer is designed to be identical to the feedforward layer it replaces. However, determining the optimal placement of the MoE layer is more of an art than a precise technical process. For the ViT-based MoE models (V-MoE), the MoE layers are positioned in the last two even-numbered blocks, as suggested in (Riquelme et al., 2021; Videau et al., 2024). In the case of SwinV1-based and SwinV2-based MoE models, the placement strategy varies depending on the dataset. For CIFAR-100 and Tiny-ImageNet, we adopt the approach from (Hwang et al., 2023), placing MoE layers in alternate layers within the last two stages. For ImageNet-1K, following (Videau et al., 2024), MoE layers are inserted exclusively into the last block of each of the last two stages for both SwinV1-based and SwinV2-based MoE models. We set the capacity ratio of MoE models to 1.25 for CIFAR-100 and Tiny-ImageNet, and to 4 for ImageNet-1K. Following (Hwang et al., 2023), a resolution of $192\times 192$ is used for SwinV2-based models during training on ImageNet-1K. All MoE layers consist of 32 experts with top-1 routing. We conduct evaluations in both training from scratch (i.e., pretraining) and fine-tuning scenarios. For fine-tuning on ImageNet-1K, MoE models perform sparse upcycling to initialize experts from dense checkpoints pretrained on ImageNet-22K (Komatsuzaki et al., 2023). During pretraining, all models are run for 150 epochs, and for fine-tuning, they are trained for 30 epochs, as in (Liu et al., 2022a, 2021). We incorporate batch prioritized routing and load balancing losses as in (Riquelme et al., 2021). MoGE extends the MoE by introducing our group sparse regularization, while all other aspects remain unchanged. Specifically, MoGE reshapes the representation $z\in\mathbb{R}^{32}$ into a 2D matrix $Z\in\mathbb{R}^{4\times 8}$ and applies a $3\times 3$ Gaussian lowpass filter. For more training details, please refer to Appendix B.
(a) Tiny-ImageNet
(b) ImageNet-1K (fine-tuning)
[…]
Figure 4. Training loss and top-1 accuracy curves of dense, MoE, and MoGE models. (a) illustrates training from scratch on Tiny-ImageNet, while (b) showcases fine-tuning on ImageNet-1K.
@@
Results. Table 2 presents the top-1 accuracy (Acc@1) and top-5 accuracy (Acc@5) achieved by various models across different datasets. The larger the value, the better the result, with the best results highlighted in bold. Notably, MoGE consistently outperforms MoE across all datasets in both top-1 and top-5 accuracies. When trained from scratch on ImageNet-1K, MoGE is inferior to its dense counterpart. This finding aligns with existing studies (Huang et al., 2023; Xue et al., 2022; Videau et al., 2024), which indicate that MoE models directly duplicated dense models fail to surpass the performance of their dense counterparts on ImageNet-1K. However, these MoE models demonstrate superior performance on ImageNet-22K. Due to the prohibitive computational cost of training from scratch on ImageNet-22K, we instead fine-tune models that are pretrained on ImageNet-22K. We can see that MoGE achieves better top-1 accuracy than the dense model, while its top-5 accuracy remains comparable.
MoGE and MoE exhibit nearly identical total parameter count (#Params) and active parameter count per token (#Paramsact) across various datasets. Although the total parameters of both MoGE and MoE are substantially larger than those of dense models, their active parameter counts per token remain nearly equivalent. This highlights the advantages of the sparsely activated MoE architecture in scaling up the number of parameters.
Fig. 4 presents parts of the training loss and the top-1 accuracy curves, with additional visualizations provided in Appendix A. Compared to MoE, MoGE incorporates a regularization constraint, which results in its training loss not consistently being lower than that of MoE. However, its top-1 accuracy is significantly higher than that of MoE.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Multimedia Asia 2025; December 09–12, 2025; Kuala Lumpur, Malaysiaisbn: 978-1-4503-XXXX-X/2018/06submissionid: 27ccs: Computing methodologies Neural networks
@@ 1. Introduction
Despite their advantages, the vanilla MoE models suffer from insufficient diversity and specialization among experts (Sun et al., 2024; Wang et al., 2024; Feng et al., 2025). In each MoE layer, the gating network determines token-to-expert assignments and should route similar tokens to the same experts to promote diversity and specialization. While load balancing losses (Shazeer et al., 2017; Lepikhin et al., 2021) encourage even token distribution, they do not ensure that similar tokens activate similar experts. As the number of experts grows, promoting diversity and specialization remains a challenge for MoE models.
@@ 2.2. Sparse Representation for MoE
2.3. Mixture of Group Experts
Table 1. Comparison of the average image Euclidean distances in $Z$ between the original images from the Fashion-MNIST test set and their corresponding variations under different transformations. $\downarrow$ indicates that smaller values are better.
Type Rotation $\downarrow$ Scaling $\downarrow$ Translation $\downarrow$ Shear $\downarrow$
@@ 2.3. Mixture of Group Experts
We first introduce the computation process of our group sparse regularization for $z$ and then provide an analysis of the underlying reasoning. Given a vector $z\in\mathbb{R}^{n}$, we begin by reshaping it into an approximately square matrix $Z\in\mathbb{R}^{r\times c}$, where $n=r\times c$ and $r$ is chosen to be as close to $c$ as possible. Next, we compute the elementwise square of $Z$ to obtain $\tilde{Z}$, i.e., $\tilde{Z}=Z\odot Z$, where $\odot$ denotes the elementwise product. Subsequently, we generate a rotationally symmetric Gaussian lowpass filter $F_{\sigma}$ of size $h\times h$ with a standard deviation $\sigma$. We then convolve $\tilde{Z}$ with the Gaussian filter $F_{\sigma}$ to produce $\tilde{P}$, i.e., $\tilde{P}=F_{\sigma}*\tilde{Z}$, where * denotes the 2D convolution. Finally, we compute the square root of each element in $\tilde{P}$ to obtain $P$. The group sparse regularization $R(z)$ is defined as the sum of all elements in $P$. We summarize the entire computation process of our group sparse regularization in Algorithm 1.
@@
where $\sigma_{0}>0$ and $\sigma_{\min}>0$ are constants satisfying $\sigma_{0}\geq\sigma_{\min}$, $\gamma\geq 0$, $t$ represents the current iteration number, $\gamma$ is the exponent, and $T$ is the total number of iterations. The curves of $\sigma$ with varying $\gamma$ over iterations are illustrated in Appendix A. Finding the optimal scheduler for $\sigma$ is a challenging task. We will explore optimal scheduler for $\sigma$ in the future.
@@
where $\lambda>0$ balances the original loss function and the regularization, and is set as described in Appendix B. Our group sparse regularization requires computing a 2D convolution with a Gaussian lowpass filter (see Algorithm 1). In practice, we empirically use valid convolution to avoid padding operations on $\tilde{Z}$. Additionally, we need to predefine the filter size $h$ and the standard deviation $\sigma$. We conduct ablation experiments in subection 3.3.
@@ 2.4. Investigating Invariant Representations
We explore how MoGE learns invariant representations by employing a simplified model with a single MoE layer for image classification on the Fashion-MNIST dataset (Xiao et al., 2017). The model consists of 400 homogeneous experts with top-1 routing, where each image is treated as a token. Each expert is a two-layer feedforward network, featuring 784 neurons in both the input and output layers, and 64 neurons in the hidden layer. The output of the MoE layer is fed directly into a single-layer feedforward classifier for categorization.
We train MoE and MoGE models from scratch for 150 epochs, setting $h=3$, $\sigma=2$, and $\lambda=$ 4e-3 for MoGE. The top-1 test accuracies achieved by MoE and MoGE are $41.70\%$ and $44.74\%$, respectively. Due to the large number of experts and the relatively small size of the training set, these test accuracies are not particularly high. However, our primary focus here is on the relative performance of the two models. Fig. 2 illustrates the average $Z$ values in Algorithm 1 for both MoE and MoGE across each category in the Fashion-MNIST test set. Notably, the average elements of $Z$ in MoGE exhibit a clear aggregation effect throughout the training process. In contrast, this aggregation effect is not observed in MoE. This indicates that the experts activated for each category in MoGE exhibit spatial similarity, which helps enhance expert diversity and specialization.
To validate that MoGE effectively learns invariant representations, we apply various transformations to the images in the Fashion-MNIST test set and evaluate the image Euclidean distances (IMED) (Wang et al., 2005) in $Z$ between the original and the transformed images. IMED accounts for spatial correlations among pixels, avoiding the perceptual pitfalls of the standard Euclidean distance, and is widely used for 2D image comparison. As shown in Table 1, we use the default function RandomAffine in PyTorch to implement four transformations: rotation, scaling, translation, and shear. The results reported in Table 1 demonstrate that following these transformations, the average IMED values in $Z$ for MoGE are consistently smaller than those for MoE. This highlights the significant advantage of MoGE in learning more invariant representations.
@@ 3. Experiments
In this section, we evaluate the superiority of MoGE on image classification and language modeling tasks, followed by ablation studies.
Table 2. Comparison of MoE and our MoGE for pretraining and fine-tuning on image classification.
Dataset Model $n$ $k$ Acc@1 Acc@5
CIFAR-100 V-MoE-S 32 1 74.03 91.49
V-MoGE-S 32 1 74.67 91.55
SwinV2-MoE-T 32 1 77.35 92.41
SwinV2-MoGE-T 32 1 77.75 92.61
Tiny-ImageNet V-MoE-S 32 1 55.38 76.66
V-MoGE-S 32 1 56.15 77.24
SwinV2-MoE-T 32 1 59.01 78.97
SwinV2-MoGE-T 32 1 60.19 79.50
ImageNet-1K V-MoE-S 32 1 76.95 92.95
V-MoGE-S 32 1 77.24 93.18
SwinV2-MoE-T 32 1 79.14 94.00
SwinV2-MoGE-T 32 1 79.97 94.94
V-MoE-S 32 1 81.19 96.07
ImageNet-1K V-MoGE-S 32 1 81.50 96.19
(fine-tuning) SwinV1-MoE-T 32 1 80.81 95.86
SwinV1-MoGE-T 32 1 81.06 95.92
@@ 3.1. Image Classification
Setup. Experiments are conducted on three widely used datasets: CIFAR-100 (Krizhevsky and Hinton, 2009), Tiny-ImageNet (Le and Yang, 2015), and ImageNet-1K (Russakovsky et al., 2015). For CIFAR-100 and Tiny-ImageNet, we apply data augmentation techniques as described in (Lee et al., 2021), while adhering to standard practices (Liu et al., 2022a, 2021; Wang et al., 2021; Liu et al., 2022b) for ImageNet-1K. We include three popular Transformer models for comparison: Vision Transformer (ViT) (Dosovitskiy et al., 2020), Swin Transformer V1 (SwinV1) (Liu et al., 2021), and Swin Transformer V2 (SwinV2) (Liu et al., 2022a).
@@
Consistent with prior studies, the architecture of each expert within the MoE layer is designed to be identical to the feedforward layer it replaces. However, determining the optimal placement of the MoE layer is more of an art than a precise technical process. For the ViT-based MoE models (V-MoE), the MoE layers are positioned in the last two even-numbered blocks, as suggested in (Riquelme et al., 2021; Videau et al., 2024). In the case of SwinV1-based and SwinV2-based MoE models, the placement strategy varies depending on the dataset. For CIFAR-100 and Tiny-ImageNet, we adopt the approach from (Hwang et al., 2023), placing MoE layers in alternate layers within the last two stages. For ImageNet-1K, following (Videau et al., 2024), MoE layers are inserted exclusively into the last block of each of the last two stages for both SwinV1-based and SwinV2-based MoE models. We conduct evaluations in both training from scratch (i.e., pretraining) and fine-tuning scenarios. For fine-tuning on ImageNet-1K, MoE models perform sparse upcycling to initialize experts from dense checkpoints pretrained on ImageNet-22K (Komatsuzaki et al., 2023). During pretraining, all models are run for 150 epochs, and for fine-tuning, they are trained for 30 epochs, as in (Liu et al., 2022a, 2021). We incorporate batch prioritized routing and load balancing losses as in (Riquelme et al., 2021). MoGE extends the MoE by introducing our group sparse regularization, while all other aspects remain unchanged. Specifically, MoGE reshapes the representation $z\in\mathbb{R}^{32}$ into a 2D matrix $Z\in\mathbb{R}^{4\times 8}$ and applies a $3\times 3$ Gaussian lowpass filter. For more training details, please refer to Appendix B.
@@
Results. Table 2 presents the top-1 accuracy (Acc@1) and top-5 accuracy (Acc@5) achieved by various models across different datasets. The larger the value, the better the result, with the best results highlighted in bold. Notably, MoGE consistently outperforms MoE across all datasets in both top-1 and top-5 accuracies. Due to the prohibitive computational cost of pretraining on ImageNet-22K, we instead fine-tune models that are pretrained on ImageNet-22K. We can see that MoGE achieves better top-1 and top-5 accuracy than MoE. The curves of training loss and top-1 accuracy for each model are provided in Appendix A.